1 Received Dec. Received $26,000 cash from Christian in exchange for commorn 2 5 9 stock. Received $3,200 cash from customers for services performed. Paid $450 cash for office supplies. Performed services for a customer and billed the customer for services rendered, $2,400. 10 Received $500 invoice for utilities due in two weeks. 15 Paid for advertising in the local paper, $275. 20 Paid utility invoice received on December 10 25 Collected cash in full from customer billed on December 9. 28 Paid rent for the month, $1,400 28 Paid $1,250 to assistant for wages. 30 Received $1,600 cash from customers for services performed. 31 Cash dividends of $5,500 were paid to stockholders.Analyze the effects of the transactions on the accounting equation of Tinnemeyer Gymnastics. Analyze the events chronologically, one transaction at a time, beginning with the transaction on the 1st. For each transaction that follows the transaction on the 1st, calculate the balance in each account after anal its effect on the accounting equation. (Complete only the necessary answer boxes for your transaction lines. [Do not enter any zeros for your transaction lines.] Carry down all balances to the Bal. line, including balance accounts, entering a .0 for any zero balances. Enter a decrease in an account with a minus sign or parentheses. Abbreviations used: AP Accounts Payable; A/R = Accounts Receivable: Adv. = Advert Com.-Common; Contr. = Contributed; Div.-Dividends; Exp.-Expense; Liab. = Liabilities: Off. Sup.-Office Supplies; Sal.-Salaries; Ser. Rev. = Service Revenue: Util. = Utility.) Assets Equity Contr Capital Retained Earnings Cash AR +Off.AP Com Div.+Serv.RentUtil.- SalAdv Rev ExpExp Exp Exp Sup Stock 12/1 +26,000
Answer #1
ASSETS Liability + Equity
Contributed Capital + Retained eArning
Cash + A/R + Office Supplies = A/P + Common Stock - Dividend + Service Revenue - Rent Expenses - Utility Expenses - Salaries Expenses - Advert. Expenses
Balance + + = + - + - - - -
12/1 $        26,000 + + = + 26000 - + - - - -
12/2 $          3,200 + + = + - + 3200 - - - -
12/5 $            -450 + + 450 = + - + - - - -
12/9 + $   2,400 + = + - + $         2,400 - - - -
12/10 + + = $       500 + - + - - 500 - -
12/15 $            -275 + + = + - + - - - - 275
12/20 $            -500 + + = $      -500 + - + - - - -
12/25 $          2,400 + $ -2,400 + = + - + - - - -
12/28 $        -1,250 + = + - + - - - 1250 -
12/30 $          1,600 + = + - + 1600 - - - -
12/31 $        -5,500 + = + - 5500 + - - - -
Closing Balance $        25,225 + $          -   + $         450 = $           -   + $                 26,000 - $        5,500 + $         7,200 - $           -   - $        500 - $    1,250 - $        275